OpenStudy (anonymous):

The perimeter of a rectangular field is 110 feet. The length is 7 feet more than the widith. Find the dimensions. System of Equations Applications! Please help

OpenStudy (anonymous):

perimeter of field = 110 ft width = x length = x+7 perimeter = 2x + 2(x+7) = 110 therefore when you do the above equation you come down to 4x + 14 = 110 minus 14 from both sides you get 96 96 divided by 4 equals 24 so that's the width and add 7 to 24 and you get the length which is 31 so dimensions are 31 by 24 hope this helps! Cheers! ^.^ Sophia